OpenStudy (roadjester):

integrate ln(x^2-x+2)

OpenStudy (roadjester):

\[\int\limits_{?}^{?}\ln (x^{2}-x+2)dx\]

OpenStudy (anonymous):

integral ln(2-x+x^2) dx Integrate by parts, integral f dg = f g- integral g df, where f = ln(x^2-x+2), dg = dx, df = (2 x-1)/(x^2-x+2) dx, g = x: = x ln(x^2-x+2)- integral (x (2 x-1))/(x^2-x+2) dx For the integrand (x (2 x-1))/(x^2-x+2), do long division: = x ln(x^2-x+2)- integral ((x-4)/(x^2-x+2)+2) dx Integrate the sum term by term: = x ln(x^2-x+2)- integral (x-4)/(x^2-x+2) dx- integral 2 dx Rewrite the integrand (x-4)/(x^2-x+2) as (2 x-1)/(2 (x^2-x+2))-7/(2 (x^2-x+2)): = x ln(x^2-x+2)- integral ((2 x-1)/(2 (x^2-x+2))-7/(2 (x^2-x+2))) dx- integral 2 dx Integrate the sum term by term and factor out constants: = x ln(x^2-x+2)+7/2 integral 1/(x^2-x+2) dx-1/2 integral (2 x-1)/(x^2-x+2) dx- integral 2 dx For the integrand (2 x-1)/(x^2-x+2), substitute u = x^2-x+2 and du = 2 x-1 dx: = -1/2 integral 1/u du+x ln(x^2-x+2)+7/2 integral 1/(x^2-x+2) dx- integral 2 dx For the integrand 1/(x^2-x+2), complete the square: = -1/2 integral 1/u du+x ln(x^2-x+2)- integral 2 dx+7/2 integral 1/((x-1/2)^2+7/4) dx For the integrand 1/((x-1/2)^2+7/4), substitute s = x-1/2 and ds = dx: = 7/2 integral 1/(s^2+7/4) ds-1/2 integral 1/u du+x ln(x^2-x+2)- integral 2 dx The integral of 1/(s^2+7/4) is (2 tan^(-1)((2 s)/sqrt(7)))/sqrt(7): = sqrt(7) tan^(-1)((2 s)/sqrt(7))-1/2 integral 1/u du+x ln(x^2-x+2)- integral 2 dx The integral of 1/u is ln(u): = sqrt(7) tan^(-1)((2 s)/sqrt(7))-(ln(u))/2+x ln(x^2-x+2)- integral 2 dx The integral of 2 is 2 x: = sqrt(7) tan^(-1)((2 s)/sqrt(7))-(ln(u))/2+x ln(x^2-x+2)-2 x+constant Substitute back for s = x-1/2: = -(ln(u))/2+x ln(x^2-x+2)-2 x+sqrt(7) tan^(-1)((2 x-1)/sqrt(7))+constant Substitute back for u = x^2-x+2: = x ln(x^2-x+2)-1/2 ln(x^2-x+2)-2 x+sqrt(7) tan^(-1)((2 x-1)/sqrt(7))+constant Which is equal to: = (x-1/2) ln(x^2-x+2)-2 x+sqrt(7) tan^(-1)((2 x-1)/sqrt(7))+constant

OpenStudy (roadjester):

@phystmg ok, now can you do that using the "SIGMA Equation" function?

OpenStudy (kinggeorge):

If you want to do a definite integral with \(\sum\) notation on an interval \([a, b]\), it would look like\[\left(b-a \over n\right)\sum_0^{n-1} {\ln(\left(a+{b-a \over n}\right)^2-\left(a+{b-a \over n}\right)+1) }\]Where a greater \(n\) corresponds to greater accuracy.

OpenStudy (roadjester):

Sorry, when I said sigma, I meant that blue button that says "Equation" and has a sigma.

OpenStudy (roadjester):

I can't translate all that wording. It's confusing.

OpenStudy (kinggeorge):

Oh well. It looks to me like phystmg just copy pasted http://www.wolframalpha.com/input/?i=ln%28x%5E2-x%2B2%29 Scroll down to where it comes up with the indefinite integral and click on show steps. It's exactly what phystmg did.

OpenStudy (roadjester):

You've got to be kidding...although useful that's just lame. If you're going to explain something, you should at least know what you're explaining.

OpenStudy (kinggeorge):

Sometimes it's just too much of a pain to write out everything yourself. And btw, I had the wrong link. It's actually http://www.wolframalpha.com/input/?i=ln%28x%5E2-x%2B1%29 The first one had \(x^2-x+2\) instead \(x^2-x+1\)

OpenStudy (roadjester):

Don't worry, you got it right the first time. It was a 2, not 1. I do appreciate the link though.

OpenStudy (kinggeorge):

I'm getting tired :( I'm not sure if there's a faster/easier way to do this. WolframAlpha generally does integrals in the worst way possible.

OpenStudy (roadjester):

lol